OpenStudy (anonymous):

Help with Stats? I will fan and medal... For a χ2 test of independence, a researcher computes a χ2 value of 22.5. The two-way table from which this value was computed had 5 rows and 4 columns. What’s the P-value for this statistic? A. P > .25 B. .025

OpenStudy (anonymous):

Degrees of freedom = (number of rows - 1) * (number of columns - 1) = 4*3 = 12 Use this and the given \(\chi^2\) value to find the p-value.

OpenStudy (amistre64):

inst a chi square always a right tail test?

OpenStudy (anonymous):

I think usually, yes, but nothing is stopping you from using a two-tailed test. I remember having used a two-tail test when estimating variance. http://stats.stackexchange.com/a/84329

OpenStudy (amistre64):

yeah, i recall that while browsing in some old stat books. wasnt real sure when they would apply tho

OpenStudy (anonymous):

Well, at any rate, I don't think this experiment is necessarily two-tailed. http://www.psychstat.missouristate.edu/introbook/sbk28m.htm

OpenStudy (anonymous):

@SithsAndGiggles and @amistre64 -- I am getting B, do you guys agree?

OpenStudy (amistre64):

id have to go home and review to be sure :)

OpenStudy (anonymous):

Okay. If you have time I would appreciate the help, if not, don't worry.

OpenStudy (anonymous):

According to this table: http://sites.stat.psu.edu/~mga/401/tables/Chi-square-table.pdf you should get a \(p\) value between \(\chi^2_{0.025}\) and \(\chi^2_{0.05}\), which are cutoff values that are equivalent to 0.025 and 0.05, respectively.

OpenStudy (amistre64):

the book i reviewed last night said that chi-square test for independence is always right tailed. It is the confidence interval for the standard deviation that got into a 2 tailed setup.

OpenStudy (amistre64):

and yeah, its making sense now :) the table doesnt give a solid Pvalue. the actual Pvalue rests within an interval, technology would need to be used to aproximate it better.